OpenStudy (anonymous):

The manager of a furniture factory finds that it costs $2400 to manufacture 90 chairs in one day and $4800 to produce 290 chairs in one day. (a) Express the cost C (in dollars) as a function of the number of chairs x produced, assuming that it is linear.

OpenStudy (anonymous):

x is chairs and y is dollars first, find the slope of this function 2400-4800 ----------- 90-290 = 12 so the slope is 12 plug in the point (90, 2400) into y-y=m(x-x) y-2400=12(x-90) y-2400=12x-1080 y=12x+1320 i THINK that's right...

OpenStudy (anonymous):

wow you made it seem so simple, thanks

OpenStudy (anonymous):

haha no prob thanks for the metal (;